In triangle ABC, a=45 degree and angle C=70 degree. Find the side opposite angle A if AB=40 cm.
Answer by Alan3354(69443)

In triangle ABC, A=45 degree and angle C=70 degree. Find the side opposite angle A if AB=40 cm.
-----------
Angle C = 180 - (45+70) = 65 degs
---
AB = side c = 40 cm
-------
a is opposite angle A
Use Law of Sines.
a/sin(A) = c/sin(C)